OpenStudy (anonymous):

You sit at the outer rim of a Ferris wheel that rotates at 2 revolutions per minute (RPM). What would your rotational speed be if you were instead clinging to a position halfway from the center to the outer rim? Wheels in rotation don’t change so no matter where everyone is it’s not going to change. So the speed would be the same.

OpenStudy (anonymous):

velocity is = w*r right? so if its going at 2revs per minute and you're further out you travel a longer distance right and if you're in closer you travel a smaller distance so your velocity should be slower So v1= 2rev/min *r v2=2rev/min*r/2 v2= 1/2v1

OpenStudy (anonymous):

Hopefully that makes sense.

OpenStudy (anonymous):

thank you